### v4

# Generative Adversarial Networks for Image Super-Resolution: A Survey

Abstract.
Single image super-resolution (SISR) has played an important role in the field of image processing. Recent generative adversarial networks (GANs) can achieve excellent results on low-resolution images with small samples. However, there are little literatures summarizing different GANs in SISR. In this paper, we conduct a comparative study of GANs from different perspectives. We first take a look at developments of GANs. Second, we present popular architectures for GANs in big and small samples for image applications. Then, we analyze motivations, implementations and differences of GANs based optimization methods and discriminative learning for image super-resolution in terms of supervised, semi-supervised and unsupervised manners, where these GANs are analyzed via integrating different network architectures, prior knowledge, loss functions and multiple tasks. Next, we compare performance of these popular GANs on public datasets via quantitative and qualitative analysis in SISR. Finally, we highlight challenges of GANs and potential research points for SISR.
SISR, GANs, small samples, optimization methods and discriminative learning
ccs: General and reference Surveys and overviews
1. Introduction
Single image super-resolution (SISR) is an important branch in the field of image processing (Wang et al., 2020a). It also aims to recover a high-resolution (HR) image over a low-resolution (LR) image (Yang et al., 2019), leading to in its wide applications in medical diagnosis (Isaac and Kulkarni, 2015), video surveillance (Zhang et al., 2010) and disaster relief (Zha et al., 2015) etc. For instance, in the medical field, obtaining higher-quality images can help doctors accurately detect diseases (Huang et al., 2017). Thus, studying SISR is very meaningful to academia and industry.
To address SISR problem, researchers have developed a variety of methods based on degradation models of low-level vision tasks (Zhou et al., 2014). There are three categories for SISR in general, i.e., image itself information, prior knowledge and machine learning. In the image itself information, directly amplifying resolutions of all pixels in a LR image through an interpolation way to obtain a HR image was a simple and efficient method in SISR (Park et al., 2003), i.e., nearest neighbor interpolation (Rukundo and Cao, 2012), bilinear interpolation (Li and Orchard, 2001) and bicubic interpolation (Keys, 1981), etc. It is noted that in these interpolation methods, high-frequency information is lost in the up-sampling process (Park et al., 2003). Alternatively, reconstruction-based methods were developed for SISR, according to optimization methods (Hardeep et al., 2013). That is, mapping a projection into a convex set to estimate the registration parameters can restore more details of SISR (Stark and Oskoui, 1989). Although the mentioned methods can overcome the drawbacks of image itself information methods, they still suffered the following challenges: non-unique solution, slow convergence speed and higher computational costs. To prevent this phenomenon, the priori knowledge and image itself information were integrated into a frame to find an optimal solution to improve the quality of the predicted SR images (Elad and Feuer, 1997; Irani and Peleg, 1991). Using maximum a posteriori (MAP) can regularize a loss function to obtain a maximum probability for improving the efficiency (Cheeseman et al., 1996). Besides, machine learning methods can be presented to deal with SISR, according to relation of data distribution (Sun et al., 2008). On the basis of ensuring the image SR effect, sparse-neighbor-embedding-based (SpNE) method via partition the training data set into a set of subsets to accelerate the speed of SR reconstruction (Gao et al., 2012). There are also many other SR methods (Sun et al., 2008; Yan et al., 2015) that often adopt sophisticated prior knowledge to restrict the possible solution space with an advantage of generating flexible and sharp detail. However, the performance of these methods rapidly degrades when the scale factor is increased, and these methods tend to be time-consuming (Nasrollahi and Moeslund, 2014).
To obtain a better and more efficient SR model, a variety of deep learning methods were applied to a large-scale image dataset to solve the super-resolution tasks. For instance, Dong et al. proposed a super-resolution convolutional neural network (SRCNN) based pixel mapping that used only three layers to obtain stronger learning ability than these of some popular machine learning methods on image super-resolution (Dong et al., 2015). Although the SRCNN had a good SR effect, it still faced problems in terms of shallow architecture and high complexity. To overcome challenges of shallow architectures, Kim et al. (Kim et al., 2016) designed a deep architecture by stacking some small convolutions to improve performance of image super-resolution. Tai et al. (Tai et al., 2017) relied on recursive and residual operations in a deep network to enhance learning ability of a SR model. To further improve the SR effect, Lee et al. (Lim et al., 2017) used weights to adjust residual blocks to achieve better SR performance. To extract robust information, the combination of traditional machine learning methods and deep networks can restore more detailed information for SISR (Wang et al., 2015). For instance, Wang et al. (Wang et al., 2015) embedded sparse coding method into a deep neural network to make a tradeoff between performance and efficiency in SISR. To reduce the complexity, an up-sampling operation is used in a deep layer in a deep CNN to increase the resolution of low-frequency features and produce high-quality images (Dong et al., 2016). For example, Dong et al. (Dong et al., 2016) directly exploited the given low-resolution images to train a SR model for improving training efficiency, where the SR network used a deconvolution layer to reconstructing HR images. There are also other effective SR methods. For example, Lai et al. (Lai et al., 2017) used Laplacian pyramid technique into a deep network in shared parameters to accelerate the training speed for SISR. Zhang et al. (Zhang et al., 2018a) guided a CNN by attention mechanisms to extract salient features for improving the performance and visual effects in image SISR.
Although the mentioned SR methods have obtained excellent effect in SISR, the obtained damaged images are insufficient in the real world, which limits the application of the above SR methods on real cameras. To address problem of small samples, generative adversarial nets (GANs) used generator and discriminator in a game-like manner to obtain good performance on image applications (Goodfellow et al., 2014; Wang et al., 2021c). Specifically, the generator can generate new samples, according existing samples (Gui et al., 2021). The discriminator is used to distinguish the samples from generator (Gui et al., 2021). Due to their strong learning abilities, GANs become popular image super-resolution methods (Bell-Kligler et al., 2019). However, there are few studies summarizing these GANs for SISR. Also, differing from previous work based deep learning techniques for image super-resolution, i.e., Refs.(Jiang et al., 2021; Anwar et al., 2020), we can not only refer to importance of GANs for low- and high-levels in terms of big and small samples, but also first deeply product a summary of GANs in image super-resolution, according to combination of different training ways (i.e., supervised, semi-supervised and unsupervised manners), network architectures, prior knowledge, loss functions and multiple tasks, which can makes readers easier know principle, improvements, superiority and inferiority of different GANs for image super-resolution. That is, in this paper, we conduct a comprehensive overview of 197 papers to show their performance, pros, cons, complexity, challenges and potential research points, etc. First, we show the effects of GANs for image applications. Second, we present popular architectures for GANs in big and small samples for image applications. Third, we analyze motivations, implementations and differences of GANs based optimization methods and discriminative learning for image super-resolution in terms of supervised, semi-supervised and unsupervised manners, where these GANs are worked by combining different network architectures, prior knowledge, loss functions and multiple tasks for image super-resolution. Fifth, we compare these GANs using experimental setting, quantitative analysis (i.e., PSNR, SSIM, complexity and running time) and qualitative analysis. Finally, we report on potential research points and existing challenges of GANs for image super-resolution. The overall architecture of this paper is shown in Fig. 1.
The remainder of this survey is organized as follows. Section 2 resents the developments of GANs. Section 3 gives a brief introduction of basic GANs for image processing tasks. Section 4 focuses on introduction of existing GANs via three ways on SISR. Section 5 compares performance of mentioned GANs from Section 4 for SISR. Section 6 offers potential directions and challenges of GANs in image super-resolution. Section 7 concludes the overview.
Figure 1. Outline of this overview. It mainly consists of basic frameworks, categories (i.e., supervised, semi-supervised and unsupervised GANs), performance comparison, challenges and potential directions.
[…]
The main structure of this paper.
2. Developments of GANs
Traditional machine learning methods prefer to use prior knowledge to improve performance of image processing applications (Sun et al., 2010). For instance, Sun et al. (Sun et al., 2010) proposed a gradient profile to restore more detailed information for improving performance of image super-resolution. Although machine learning methods based prior knowledge has fast execution speed, they have some drawbacks. First, they required manual setting parameters to achieve better performance on image tasks. Second, they required complex optimization methods to find optimized parameters. According to mentioned challenges, deep learning methods are developed (Bird et al., 2022). Deep learning methods used deep networks, i.e., CNNs to automatically learn features rather than manual setting parameters to obtain effective effects in image processing tasks, i.e., image classification (Bird et al., 2022), image inpainting (Liu et al., 2021) and image super-resolution (Wang et al., 2020a). Although these methods are effective big samples, they are limited for image tasks with small samples (Goodfellow et al., 2014).
To address problems above, GANs are presented in image processing (Goodfellow et al., 2014; Saxena and Cao, 2021). GANs consist of generator network and discriminator network. The generator network is used to generate new samples, according to given samples. The discriminator network is used to determine truth of obtained new samples. When generator and discriminator is balance, a GAN model is finished. The work process of GAN can be shown in Fig. 2, where G and D denote a generator network and discriminator network. To better understand GANs, we introduce several basic GANs as follows.
Figure 2. Architecture of generative adversarial network (GAN).
[…]
The original structure of GAN.
To obtain more realistic effects, conditional information is fused into a GAN (CGAN) to randomly generate images, which are closer to real images (Mirza and Osindero, 2014). CGAN improves GAN to obtain more robust data, which has an important reference value to GANs for computer vision applications. Subsequently, increasing the depth of GAN instead of the original multilayer perceptron in a CNN to improve expressive ability of GAN is developed for complex vision tasks (Radford et al., 2015). To mine more useful information, the bidirectional generative adversarial network (BiGAN) used dual encoders to collaborate a generator and discriminator to obtain richer information for improving performance in anomaly detection, which is shown in Fig. 3 (Donahue et al., 2016). In Fig. 3, x denotes a feature vector, E is an encoder and y expresses an image from discriminator.
Figure 3. Architecture of bidirectional generative adversarial network (BiGAN).
[…]
The structure of BiGAN.
It is known that pretrained operations can be used to accelerate the training speed of CNNs for image recognition (Hinterstoisser et al., 2018). This idea can be treated as an energy drive. Inspired by that, Zhao et al. proposed an energy-based generative adversarial network (EBGAN) by using a pretraining operation into a discriminator to improve the performance in image recognition (Zhao et al., 2016). To keep consistency of obtained features with original images, cycle-consistent adversarial network (CycleGAN) relies on a cyclic architecture to achieve an excellent style transfer effect (Zhu et al., 2017) as illustrated in Fig. 4.
Figure 4. Architecture of cycle-consistent adversarial network (CycleGAN).
[…]
The structure of CycleGAN.
Although pretrained operations are useful for training efficiency of network models, they may suffer from mode collapse. To address this problem, Wasserstein GAN (WGAN) used weight clipping to enhance importance of Lipschitz constraint to improve the stability of training a GAN (Arjovsky et al., 2017). WGAN used weight clipping to perform well. However, it is easier to cause gradient vanishing or gradient exploding (Gulrajani et al., 2017). To resolve this issue, WGAN used a gradient penalty (treated as WGAN-GP) to break the limitation of Lipschitz for pursuing good performance in computer vision applications (Brock et al., 2018). To further improve results of image generation, GAN enlarged batch size and used truncation trick as well as BIGGAN can make a tradeoff between variety and fidelity (Brock et al., 2018). To better obtained features of different parts of an image (i.e., freckles and hair), style-based GAN (StyleGAN) uses feature decoupling to control different features and finish style transfer for image generation (Karras et al., 2019). The architecture of StyleGAN and its generator are shown in Fig. 5 and Fig. 6.
Figure 5. Architecture of StyleGAN.
[…]
The structure of StyleGAN.
Figure 6. The structure of generator in the StyleGAN.
[…]
The structure of generator in the StyleGAN.
In recent years, GANs with good performance have been applied in the fields of image processing, natural language processing (NLP) and video processing. Also, there are other variants based on GANs for multimedia applications, such as Laplacian pyramid of GAN (LAPGAN) (Denton et al., 2015), coupled GAN (CoupleGAN) (Liu and Tuzel, 2016), self-attention GAN (SAGAN) (Zhang et al., 2019a), loss-sensitive GAN (LSGAN) (Qi, 2020). These methods emphasize how to generate high-quality images through various sampling mechanisms. However, researchers focused applications of GANs from 2019, i.e., FUNIT (Liu et al., 2019a), SPADE (Park et al., 2019) and U-GAT-IT (Kim et al., 2019). Illustrations of more GANs are shown in Table 1.
Table 1. Introduction of many GANs.
Models Methods Applications Key words
GAN (Goodfellow et al., 2014) GAN Image generation GAN in a semi-supervised way for image generation
DICCGAN (Bird et al., 2022) CGAN Image classification Conditional GAN for image classification
PD-GAN (Liu et al., 2021) GAN Image inpainting GAN for image inpainting and image restoration
CGAN (Mirza and Osindero, 2014) GAN Image generation GAN in a supervised way for image generation
DCGAN (Radford et al., 2015) GAN Image generation GAN in an unsupervised way for image generation
BiGAN (Donahue et al., 2016) GAN Image generation GAN with encoder in an unsupervised way for image generation
EBGAN (Zhao et al., 2016) GAN Image generation and training nets GAN based energy for image generation
CycleGAN (Zhu et al., 2017) GAN Image generation GAN with cycle-consistent for image generation
WGAN-GP (Gulrajani et al., 2017) GAN Image generation GAN with gradient penalty for image generation
BIGGAN (Brock et al., 2018) GAN Image super-resolution GAN with big channels of image super-resolution
StyleGAN (Karras et al., 2019) GAN Image generation GAN with stochastic variation for image generation
LAPGAN (Denton et al., 2015) CGAN Image super-resolution GAN with Laplacian pyramid for image super-resolution
CoupleGAN (Liu and Tuzel, 2016) GAN Image generation GAN for both up-sampling and image generation
SAGAN (Zhang et al., 2019a) GAN Image generation Unsupervised GAN with self-attention for image generation
FUNIT (Liu et al., 2019a) GAN Image translation GAN in an unsupervised way for image-to-image translation
SPADE (Park et al., 2019) GAN Image generation GAN with spatially-adaptive normalization for image generation
U-GAT-IT (Kim et al., 2019) GAN Image translation GAN with attention in an unsupervised way for image-to-image translation
3. Popular GANs for image applications
According to mentioned illustrations, it is known that variants of GANs based on properties of vision tasks are developed in Section 2. To further know GANs, we show different GANs on training data, i.e., big samples and small samples for different high- and low-level computer vision tasks as shown in Fig. 7.
Figure 7. Frame of popular GANs for image applications.
[…]
The structure of the Section 3.
3.1. GANs on big samples for image applications
3.1.1. GANs on big samples for image generation
Good performance of image generation depends on rich samples. Inspired by that, GANs are improved for image generation (Gui et al., 2021). That is, GANs use generator to produce more samples from high-dimensional data to cooperate discriminator for promoting results of image generation. For instance, boundary equilibrium generative adversarial networks (BEGAN) used obtained loss from Wasserstein to match loss of auto-encoder in the discriminator and achieve a balance between a generator and discriminator, which can obtain more texture information than that of common GANs in image generation (Berthelot et al., 2017). To control different parts of a face, StyleGAN decoupled different features to form a feature space for finishing transfer of texture information (Karras et al., 2019). Besides, texture synthesis is another important application of image generation (Li and Wand, 2016). For instance, Markovian GANs (MGAN) can quickly capture texture date of Markovian patches to achieve function of real-time texture synthesis (Li and Wand, 2016), where Markovian patches can be obtained Ref. (Gui et al., 2021). Periodic spatial GAN (PSGAN) (Bergmann et al., 2017) is a variant of spatial GAN (SGAN) (Jetchev et al., 2016), which can learn periodic textures of big datasets and a single image. These methods can be summarized in Table 2.
Table 2. GANs on big samples for image generation.
Models Methods Key words
StyleGAN (Karras et al., 2019) GAN GAN with stochastic variation for image generation
BEGAN (Berthelot et al., 2017) GAN GAN with upsampling for image generation
MGAN (Li and Wand, 2016) GAN GAN with Markovian for texture synthesis
PSGAN (Bergmann et al., 2017) GAN Periodic GAN for texture synthesis
SGAN (Jetchev et al., 2016) GAN GAN with spatial tensor for texture synthesis
3.1.2. GANs on big samples for object detection
Object detection has wide applications in the industry, i.e., smart transportation (Wang et al., 2021a) and medical diagnosis (Gerdprasert and Mabu, 2021), etc. However, complex environments have huge challenges for pursuing good performance of object detection methods (Zou et al., 2019). Rich data is important for object detection. Existing methods used a data-driven strategy to collect a large-scale dataset including different object examples under different conditions to obtain an object detector. However, the obtained dataset does not contain all kinds of deformed and occluded objects, which limits effects of object detection methods. To resolve the issue, GANs are used for object detection (Ehsani et al., 2018; Li et al., 2017a). Ehsani et al. used segmentation and generation in a GANs from invisible parts in the objects to overcome occluded objects (Ehsani et al., 2018). To address a challenge of small object detection on low-resolution and noisy representation, a perceptual GAN (Perceptual GAN) reduced differences of small objects and big objects to improve performance in small object detection (Li et al., 2017a). That is, its generator converted poor perceived representation from small objects to high-resolution big objects to fool a discriminator, where mentioned big objects are similar to real big objects (Li et al., 2017a). To obtain sufficient information of objects, an end-to-end multi-task generative adversarial network (SOD-MTGAN) used a generator to recover detailed information for generating high-quality images for achieving accurate detection (Bai et al., 2018). Also, a discriminator transferred classification and regression losses in a back-propagated way into a generator (Bai et al., 2018). Two operations can extract objects from backgrounds to achieve good performance in object detection [59]. More detailed information is shown in Table 3.
Table 3. GANs on big samples for object detection.
Models Methods Key words
SeGAN (Ehsani et al., 2018) GAN GAN with segmentor for object detection
Perceptual GAN (Li et al., 2017a) GAN GAN with super-resolved representation for object detection
SOD-MTGAN (Bai et al., 2018) GAN Multi-task GAN for object detection
3.2. GANs on small samples for image applications
3.2.1. GANs on small samples for image style transfer
Makeup has important applications in the real world (Yuan and Zhang, 2022). To save costs, visual makeup software is developed, leading to image style transfer (i.e., image-to-image) translation becoming a research hotspot in the field of computer vision in recent years (Gui et al., 2021). GANs are good tools for style transfer on small samples, which can be used to establish mappings between given images and object images (Gui et al., 2021). The obtained mappings are strongly related to aligned image pairs (Isola et al., 2017). However, we found that the above mappings do not match our ideal models in terms of transfer effects (Zhu et al., 2017). Motivated by that, CycleGAN used two pairs of a generator and discriminator in a cycle consistent way to learn two mappings for achieving style transfer (Zhu et al., 2017). CycleGAN had two phases in style transfer. In the first phase, an adversarial loss (Park et al., 2019) was used to ensure the quality of generated images. In the second phase, a cycle consistency loss (Zhu et al., 2017) was utilized to guarantee that predicted images to fell into the desired domains (Chang et al., 2018). CycleGAN had the following merits. It does not require paired training examples (Chang et al., 2018). And it does not require that the input image and the output image have the same low-dimensional embedding space (Zhu et al., 2017). Due to its excellent properties, many variants of CycleGAN have been conducted for many vision tasks, i.e., image style transfer (Zhu et al., 2017; Chen et al., 2021), object transfiguration (Kim and Park, 2018) and image enhancement (You et al., 2019b), etc. More GANs on small samples for image style transfer can be found in Table 4.
Table 4. GANs on small samples for image style transfer.
Models Methods Key words
RAMT-GAN (Yuan and Zhang, 2022) GAN GAN for image style transfer on makeup
CycleGAN (Zhu et al., 2017) GAN Cycle-consistent GAN for image-to-image translation
CATVGAN (Bin et al., 2020) GAN Correlation alignment GAN for image style transfer
ITCGAN (Isola et al., 2017) CGAN CGAN with U-net for image-to-image translation
ArCycleGAN (Chen et al., 2021) GAN GAN with attribute registration for image-to-image translation
URCycleGAN (Kim and Park, 2018) CycleGAN CycleGAN with U-net for image-to-image translation
ECycleGAN (You et al., 2019b) CycleGAN CycleGAN with convolutional block attention module (CBAM) for image-to-image translation
3.2.2. GANs on small samples for image inpainting
Images have played important roles in human–computer interaction in the real world (Cowie et al., 2001). However, they may be damaged when they were collected by digital cameras, which has a negative impact on high-level computer vision tasks. Thus, image inpainting had important values in the real world (Guillemot and Le Meur, 2013). Due to missing pixels, image inpainting suffered from enormous challenges (Elharrouss et al., 2020). To overcome shortcoming above, GANs are used to generate useful information to repair damaged images based on the surrounding pixels in the damaged images (Demir and Unal, 2018). For instance, GAN used a reconstruction loss, two adversarial losses and a semantic parsing loss to guarantee pixel faithfulness and local-global contents consistency for face image inpainting (Li et al., 2017b). Although this method can generate useful information, which may cause boundary artifacts, distorted structures and blurry textures inconsistent with surrounding areas (Zhang et al., 2022b; Yu et al., 2018). To resolve this issue, Zhang et al. embedded prior knowledge into a GAN to generate more detailed information for achieving good performance in image inpainting (Zhang et al., 2022b). Yu et al. exploited a contextual attention mechanism to improve a GAN for obtaining excellent visual effect in image inpainting (Yu et al., 2018). Typical GANs on small samples for image inpainting is summarized in Table 5.
Table 5. GANs on small samples for image inpainting.
Models Methods Key words
PGGAN (Demir and Unal, 2018) GAN GAN based patch for image inpainting
DE-GAN (Zhang et al., 2022b) GAN GAN with prior knowledge for face inpainting
GFC (Li et al., 2017b) GAN GAN with autoencoder for image inpainting
GIICA (Yu et al., 2018) WGAN WGAN with attention model for image inpainting
4. GANs for image super-resolutions
According to mentioned illustrations, it is clear that GANs have many important applications in image processing. Also, image super-resolution is crucial for high-level vision tasks, i.e., medical image diagnosis and weather forecast, etc. Thus, GANs in image super-resolution have important significance in the real world. However, there are few summaries about GANs for image super-resolution. Inspired by that, we show GANs for image super-resolution, according to supervised GANs, semi-supervised GANs and unsupervised GANs for image super-resolution as shown in Fig. 8. Specifically, supervised GANs in image super-resolution include supervised GANs based improved architectures, supervised GANs based prior knowledge, supervised GANs with improved loss functions and supervised GANs based multi-tasks for image super-resolution. Semi-supervised GANs for image super-resolution contain semi-supervised GANs based improved architectures, semi-supervised GANs with improved loss functions and semi-supervised GANs based multi-tasks for image super-resolution.
Unsupervised GANs for image super-resolution consists of unsupervised GANs based improved architectures, unsupervised GANs based prior knowledge, unsupervised GANs with improved loss functions and unsupervised GANs based multi-tasks in image super-resolution. More information of GANs on image super-resolution can be illustrated as follows.
4.1. Supervised GANs for image super-resolution
4.1.1. Supervised GANs based improved architectures for image super-resolution
GANs in a supervised way to train image super-resolution models are very mainstream. Also, designing GANs via improving network architectures are very novel. Thus, improved GANs in a supervised way for image super-resolution are very popular. That can improve GANs by designing novel discriminator networks, generator networks, attributes of image super-resolution task, complexity and computational costs. For example, Laplacian pyramid of adversarial networks (LAPGAN) fused a cascade of convolutional networks into Laplacian pyramid network in a coarse-to-fine way to obtain high-quality images for assisting image recognition task (Denton et al., 2015). To overcome the effect of big scales, curvature and highlight compact regions can be used to obtain a local salient map for adapting big scales in image-resolution (Mahapatra et al., 2017). More research on improving discriminators and generators is shown as follows.
In terms of designing novel and discriminators and generators, progressive growing generative adversarial networks (PGGAN or ProGAN) utilized different convolutional layers to progressively enlarge low-resolution images to improve image qualities for image recognition (Karras et al., 2017). An enhanced SRGAN (ESRGAN) used residual dense blocks into a generator without batch normalization to mine more detailed information for image super-resolution (Wang et al., 2018b). To eliminate effects of checkerboard artifacts and the unpleasing high-frequency, multi-discriminators were proposed for image super-resolution (Lee et al., 2019). That is, a perspective discriminator was used to overcome checkerboard artifacts and a gradient perspective was utilized to address unpleasing high-frequency question in image super-resolution. To improve the perceptual quality of predicted images, ESRGAN+ fused two adjacent layers in a residual learning way based on residual dense blocks in a generator to enhance memory abilities and added noise in a generator to obtain stochastic variation and obtain more details of high-resolution images (Rakotonirina and Rasoanaivo, 2020).
Restoring detailed information may generate artifacts, which can seriously affect qualities of restored images (Zhang and Ling, 2020). In terms of face image super-resolution, Zhang et al. used a supervised pixel-wise GAN (SPGAN) to obtain higher-quality face images via given low-resolution face images of multiple scale factors to remove artifacts in image super-resolution (Zhang and Ling, 2020). In terms of remote sensing image super-resolution, Gong et al. used enlighten blocks to make a deep network achieve a reliable point and used self-supervised hierarchical perceptual loss to overcome effects of artifacts in remote sensing image super-resolution (Gong et al., 2021). Dharejo et al. used Wavelet Transform (WT) characteristics into a transferred GAN to eliminate artifacts to improve quality of predicted remote sensing images (Dharejo et al., 2021). Moustafa et al. embedded squeeze-and-excitation blocks and residual blocks into a generator to obtain more high-frequency details (Moustafa and Sayed, 2021). Besides, Wasserstein distance is used to enhance the stability of training a remote sensing super-resolution model (Moustafa and Sayed, 2021). To address pseudo-textures problem, a saliency analysis is fused with a GAN to obtain a salient map that can be used to distinguish difference between a discriminator and a generator (Ma et al., 2019).
To obtain more detailed information in image super-resolution, a lot of GANs are developed (Ko and Dai, 2021). Ko et al. used Laplacian idea and edge in a GAN to obtain more useful information to improve clarities of predicted face images (Ko and Dai, 2021). Using tensor structures in a GAN can facilitate texture information for SR (Ding et al., 2019). Using multiple generators in a GAN can obtain more realistic texture details, which was useful to recover high-quality images (Zareapoor et al., 2019; Xing et al., 2021). To obtain better visual effects, a gradually GAN used gradual growing factors in a GAN to improve performance in SISR (Shamsolmoali et al., 2019).
To reduce computational costs and memory, Ma et al. used two-stage generator in a supervision way to extract more effective features of cytopathological images, which can reduce the cost of data acquisition and save cost (Ma et al., 2020). Cheng et al. designed a generator by multi-scale feature aggregation and a discriminator via a PatchGAN to reduce memory consumption for a GAN on SR (Cheng et al., 2021). Besides, distilling a generator and discriminator can accelerate the training efficiency of a GAN model for SR (Cheng et al., 2021). More supervised GANs for image super-resolution are shown in Table 6.
Figure 8. Frame of GANs for image super-resolution.
[…]
The structure of the Section 4.
Table 6. Supervised GANs for image super-resolution in section 4.1.1.
Models Methods Key words
LAPGAN (Denton et al., 2015) CGAN CGAN with Laplacian Pyramid for image super-resolution
LSMGAN (Mahapatra et al., 2017) CGAN CGAN with local saliency maps for retinal image super-resolution
PGGAN (Karras et al., 2017) GAN Progressive growing GAN for image super-resolution
ESRGAN (Wang et al., 2018b) SRGAN SRGAN with Residual-in-Residual Dense Block (RRDB) and relativistic discriminator for image super-resolution
MPDGAN (Lee et al., 2019) GAN GAN with multi-discriminators for image super-resolution
ESRGAN+ (Rakotonirina and Rasoanaivo, 2020) ESRGAN ESRGAN with Residual-in-Residual Dense Residual Block (RRDRB) for image super-resolution
SPGAN (Zhang and Ling, 2020) GAN GAN with identity-based discriminator for face image super-resolution
MLGE (Ko and Dai, 2021) LAPGAN LAPGAN with edge information for face image super-resolution
SD-GAN (Ma et al., 2019) GAN GAN for remote sensing image super-resolution
PathSRGAN (Ma et al., 2020) SRGAN SRGAN with RRDB for cytopathology image super-resolution
Enlighten-GAN (Gong et al., 2021) GAN GAN with enlighten block for remote sensing image super-resolution
TWIST-GAN (Dharejo et al., 2021) GAN GAN with wavelet transform (WT) for remote sensing image super-resolution
SCSE-GAN (Moustafa and Sayed, 2021) GAN GAN with SCSE block for image super-resolution
MFAGAN (Cheng et al., 2021) GAN GAN with multi-scale feature aggregation net for image super-resolution
TGAN (Ding et al., 2019) GAN GAN with visual tracking and attention networks for image super-resolution
DGAN (Zareapoor et al., 2019) GAN GAN with disentangled representation learning and anisotropic BRDF reconstruction for image super-resolution
DMGAN (Xing et al., 2021) GAN GAN with two same generators for image super-resolution
G-GANISR (Shamsolmoali et al., 2019) GAN GAN with gradual learning for image super-resolution
SRGAN (Ledig et al., 2017) GAN GAN with deep ResNet for image super-resolution
RaGAN (Jolicoeur-Martineau, 2018) GAN GAN with relativistic discriminator for image super-resolution
LE-GAN (Shi et al., 2021) GAN GAN with a latent encoder for realistic hyperspectral image super-resolution
NCSR (Kim and Son, 2021) GAN GAN with a noise conditional layer for image super-resolution
Beby-GAN (Li et al., 2021) GAN GAN with a region-aware adversarial learning strategy for image super-resolution
MA-GAN (Xu et al., 2021) GAN GAN with pyramidal convolution for image super-resolution
CMRI-CGAN (Xia et al., 2021a) CGAN CGAN with optical flow component for magnetic resonance image super-resolution
D-SRGAN (Demiray et al., 2021) SRGAN SRGAN for image super-resolution
LMISR-GAN (Ma et al., 2021) GAN GAN with residual channel attention block for medical image super-resolution
4.1.2. Supervised GANs based prior knowledge for image super-resolution
It is known that combination of discriminative method and optimization can make a tradeoff between efficiency and performance (Zhang et al., 2017). Guan et al. used high-resolution image to low-resolution image network and low-resolution image to high-resolution image network with nearest neighbor down-sampling method to learn detailed information and noise prior for image super-resolution (Guan et al., 2019). Chan et al. used rich and diverse priors in a given pretrained to mine latent representative information for generating realistic textures for image super-resolution (Chan et al., 2021b). Liu et al. used a gradient prior into a GAN to suppress the effect of blur kernel estimation for image super-resolution (Liu et al., 2019b).
4.1.3. Supervised GANs with improved loss functions for image super-resolution
Loss function can affect performance and efficiency of a trained SR model. Thus, we analyze the combination of GANs with different loss functions in image super-resolution (Zhang et al., 2019b). Zhang et al. trained a Ranker to obtain representation of perceptual metrics and used a rank-content loss in a GAN to improve visual effects in image super-resolution (Zhang et al., 2019b). To eliminate effect of artifacts, Zhu et al. used image quality assessment metric to implement a novel loss function to enhance the stability for image super-resolution (Zhu et al., 2020). To decrease complexity of GAN model in image super-resolution, Fuoli et al. used a Fourier space supervision loss to recover lost high-frequency information to improve predicted image quality and accelerate training efficiency in SISR (Fuoli et al., 2021). To enhance stability of a SR model, using residual blocks and a self-attention layer in a GAN enhances robustness of a trained SR model. Also, combining improved Wasserstein gradient penalty and perceptual Loss enhances stability of a SR model (Shahidi, 2021). To extract accurate features, fusing a measurement loss function into a GAN can obtain more detailed information to obtain clearer images (Shahsavari et al., 2021).
4.1.4. Supervised GANs based multi-tasks for image super-resolution
Improving image quality is important for high-level vision tasks, i.e., image recognition (Talab et al., 2019). Besides, devices often suffer from effects of multiple factors, i.e., device hardware, camera shakes and shooting distances, which results in collected images are damaged. That may include noise and low-resolution pixels. Thus, addressing the multi-tasks for GANs are very necessary (Hu et al., 2019). For instance, Adil et al. exploited SRGAN and a denoising module to obtain a clear image. Then, they used a network to learn unique representative information for identifying a person (Adil et al., 2020). In terms of image super-resolution and object detection, Wang et al. used multi-class cyclic super-resolution GAN to restore high-quality images, and used a YOLOv5 detector to finish object detection task (Wang et al., 2022). Zhang et al. used a fully connected network to implement a generator for obtaining high-definition plate images and a multi-task discriminator is used to enhance super-resolution and recognition tasks (Zhang et al., 2018b). The use of an adversarial learning was a good tool to simultaneously address text recognition and super-resolution (Xu et al., 2020).
In terms of complex damaged image restoration, GANs are good choices (Li et al., 2022). For instance, Li et al. used a multi-scale residual block and an attention mechanism in a GAN to remove noise and restore detailed information in CTA image super-resolution (Li et al., 2022). Nneji et al. improved a VGG19 to fine-tune two sub-networks with a wavelet technique to simultaneously address COVID-19 image denoising and super-resolution problems (Nneji et al., 2022). More information is shown in Table 7.
Table 7. Supervised GANs for image super-resolution in section 4.1.2 to section 4.1.4.
Models Methods Key words
SRDGAN (Guan et al., 2019) GAN GAN with GMSR for image super-resolution
GLEAN (Chan et al., 2021b) GAN GAN with pre-trained models for image super-resolution
I-SRGAN (Liu et al., 2019b) GAN GAN with infrared prior knowledge for image super-resolution on infrared image
RankSRGAN (Zhang et al., 2019b) SRGAN SRGAN with ranker for image super-resolution
GMGAN (Zhu et al., 2020) GAN GAN with a novel quality loss for image super-resolution
FSLSR (Fuoli et al., 2021) GAN GAN with fourier space losses for image super-resolution
I-WAGAN (Shahidi, 2021) GAN GAN with improved wasserstein gradient penalty and perceptual loss for image super-resolution
CESR-GAN (Shahsavari et al., 2021) GAN GAN with a feature-based measurement loss function for image super-resolution
RTSRGAN (Hu et al., 2019) SRGAN SRGAN for real time image super-resolution
MSSRGAN (Adil et al., 2020) ESRGAN ESRGAN with denoising module for image super-resolution
RSISRGAN (Wang et al., 2022) GAN GAN for image super-resolution on RSI
JPLSRGAN (Zhang et al., 2018b) GAN GAN for license plate recognition and image super-resolution
SRR-GAN (Xu et al., 2020) GAN GAN for image super-resolution on text images
MRD-GAN (Li et al., 2022) GAN GAN with attention mechanism for image super-resolution and denoising
MESRGAN+ (Nneji et al., 2022) ESRGAN ESRGAN with siamese network for image super-resolution and denoising.
RealESRGAN (Wang et al., 2021a) ESRGAN ESRGAN with pure synthetic data for blind image super-resolution
SNPE-SRGAN (Tampubolon et al., 2021) SRGAN SRGAN with SPNE for image super-resolution
SOUP-GAN (Zhang et al., 2022a) GAN GAN with 3D MRI for image super-resolution
4.2. Semi-supervised GANs for image super-resolution
4.2.1. Semi-supervised GANs based improved architectures for image super-resolution
For real problems with less data, semi-supervised techniques are developed. For instance, asking patients takes multiple CT scans with additional radiation doses to conduct paired CT images for training SR models in clinical practice is not realistic. Motivated by that, GANs in semi-supervised ways are used for image super-resolution (You et al., 2019a). For instance, by maintaining the cycle-consistency of Wasserstein distance, a mapping from noisy low-resolution images to high-resolution images was built (You et al., 2019a). Besides, combining a convolutional neural network, residual learning operations in a GAN can facilitate more detailed information for image super-resolution (You et al., 2019a). To resolve super-resolution and few labeled samples, Xia et al. used soft multi-labels to implement a semi-supervised super-resolution method for person re-identification (Xia et al., 2021b). That is, first, a GAN is used to conduct a SR model. Second, a graph convolutional network is exploited to construct relationship of local features from a person. Third, some labeled samples are used to train unlabeled samples via a graph convolutional network.
4.2.2. Semi-supervised GANs with improved loss functions and semi-supervised GANs based multi-tasks for image super-resolution
The combinations of semi-supervised GANs and loss functions are also effective in image super-resolution (Jiang et al., 2020). For example, Jiang et al. combined an adversarial loss, a cycle-consistency loss, an identity loss and a joint sparsifying transform loss into a GAN in a semi-supervised way to train a CT image super-resolution model (Jiang et al., 2020). Although this model made a significantly progress on some evaluation criteria, it was still disturbed by artifacts and noise.
In terms of multi-tasks, Nicolo et al. proposed to use a mixed adversarial Gaussian domain adaptation in a GAN in a semi-supervised way to obtain more useful information for implementing a 3D super-resolution and segmentation (Savioli et al., 2021). More information of semi-supervised GANs in image super-resolution can be illustrated in Table 8.
Table 8. Semi-supervised GANs for image super-resolution in section 4.2.
Models Methods Key words
GAN-CIRCLE (You et al., 2019a) GAN GAN with cycle-consistency of Wasserstein distance in a semi-supervised way for noisy image super-resolution
MSSR (Xia et al., 2021b) GAN GAN with soft multi-labels in a semi-supervised way for image super-resolution
CTGAN (Jiang et al., 2020) GAN GAN with four losses in a semi-supervised way for image super-resolution
Gemini-GAN (Savioli et al., 2021) GAN GAN with mixed adversarial Gaussian domain adaptation in a semi-supervised way for 3D super-resolution and segmentation
4.3. Unsupervised GANs for image super-resolution
Collected images in the real world have less pairs. To address this phenomenon, unsupervised GANs are presented (Yuan et al., 2018). It can be divided into four types, i.e., improved architectures, prior knowledge, loss functions and multi-tasks in GANs in unsupervised ways for image super-resolution as follows.
4.3.1. Unsupervised GANs based improved architectures for image super-resolution
CycleGANs have obtained success in unsupervised ways in image-to-image translation applications (Zhu et al., 2017). Accordingly, the CycleGANs are extended into SISR to address unpair images (i.e., low-resolution and high-resolution) in the datasets in the real world (Yuan et al., 2018). Yuan et al. used a CycleGAN for blind super-resolution over the following phases (Yuan et al., 2018). The first phase removed noise from noisy and low-resolution images. The second phase resorted to an up-sampled operation in a pre-trained deep network to enhance the obtained low-resolution images. The third phase used a fine-tune mechanism for a GAN to obtain high-resolution images. To address blind super-resolution, bidirectional structural consistency was used into a GAN in an unsupervised way to train a blind SR model and construct high-quality images (Zhao et al., 2018). Alternatively, Zhang et al. exploited multiple GANs as basis components to implement an improved CycleGAN for train an unsupervised SR model (Zhang et al., 2019c). To eliminate checkerboard artifacts, an upsampling module containing a bilinear interpolation and a transposed convolution was used in an unsupervised CycleGAN to improve visual effects of restored images in the real world (Kim et al., 2020).
There are also other popular methods that use GANs in unsupervised ways for image super-resolution (Prajapati et al., 2020). To improve the learning ability of a SR model in the real world, it combines an unsupervised learning and a mean opinion score in a GAN to improve perceptual quality in the real-world image super-resolution (Prajapati et al., 2020). To recover more natural image characteristics, Lugmayr et al. combined unsupervised and supervised ways for blind image super-resolution (Lugmayr et al., 2019). The first step learned to invert the effects of bicubic down sampling operation in a GAN in an unsupervised way to extract useful information from natural images (Lugmayr et al., 2019). To generate image pairs in the real world, the second step used a pixel-wise network in a supervised way to obtain high-resolution images (Lugmayr et al., 2019). To break fixed downscaling kernel, Sefi et al. used KernelGAN (Bell-Kligler et al., 2019) and Internal-GAN (Shocher et al., 2018) to obtain an internal distribution of patches in the blind image super-resolution. To accelerate the training speed, a guidance module was used in a GAN to quickly seek a correct mapping from a low-resolution domain to a high-resolution domain in unpaired image super-resolution (Lian et al., 2019). To improve the accuracy of medical diagnosis, Song et al. used dual GANs in a self-supervised way to mine high dimensional information for PET image super-resolution (Mahapatra et al., 2017). Besides, other SR methods can have an important reference value for unsupervised GANs with for image super-resolution. For example, Wang et al. used an unsupervised method to translate real low-resolution images to real low-resolution images (Wang et al., 2020b). Chen et al. resorted to a supervised super-resolution method to convert obtained real low-resolution images into real high-resolution images (Chen et al., 2020). More information of mentioned unsupervised GANs for image super-resolution can be shown in Table 9 as follows.
Table 9. Unsupervised GANs based improved architectures for image super-resolution.
Models Methods Key words
CinCGAN (Yuan et al., 2018) GAN Unsupervised GAN for image super-resolution
DNSR (Zhao et al., 2018) GAN Unsupervised GAN with bidirectional structural consistency for blind image super-resolution
MCinCGAN (Zhang et al., 2019c) CycleGAN Unsupervised GAN for image super-resolution
RWSR-CycleGAN (Kim et al., 2020) CycleGAN Unsupervised GAN for image super-resolution
USISResNet (Prajapati et al., 2020) GAN Unsupervised GAN with USISResNet for image super-resolution
ULRWSR (Lugmayr et al., 2019) GAN Unsupervised GAN with pixel wise supervision for image super-resolution
KernelGAN (Bell-Kligler et al., 2019) Internal-GAN Unsupervised GAN for blind image super-resolution
InGAN (Shocher et al., 2018) GAN Unsupervised GAN for image super-resolution
FG-SRGAN (Lian et al., 2019) SRGAN Unsupervised GAN with a guided block for image super-resolution
PETSRGAN (Mahapatra et al., 2017) GAN Unsupervised GAN with a self-supervised way for PET image super-resolution
TrGAN (Wang et al., 2020b) GAN Unsupervised GAN for image synthesis and super-resolution
CycleSR (Chen et al., 2020) GAN Unsupervised GAN with an indirect supervised path for image super-resolution
UGAN-Circle (Guha et al., 2021) GAN-Circle Unsupervised GAN-Circle for image super-resolution on CT images
4.3.2. Unsupervised GANs based prior knowledge for image super-resolution
Combining unsupervised GANs and prior knowledge in unsupervised GANs can better address unpair image super-resolution (Lin et al., 2018). Lin et al. combined data error, a regular term and an adversarial loss to guarantee consistency of local-global content and pixel faithfulness in a GAN in an unsupervised way to train an image super-resolution model (Lin et al., 2018). To better support medical diagnosis, Das et al. combined adversarial learning in a GAN, cycle consistency and prior knowledge, i.e., identity mapping prior to facilitate more useful information i.e., spatial correlation, color and texture information for obtaining cleaner high-quality images (Das et al., 2020). In terms of remoting sensing super-resolution, a random noise is used in a GAN to reconstruct satellite images (Wang et al., 2021b). Then, authors conducted image prior by transforming the reference image into a latent space (Wang et al., 2021b). Finally, they updated the noise and latent space to transfer obtained structure information and texture information for improving resolution of remote sensing images (Wang et al., 2021b).
4.3.3. Unsupervised GANs with improved loss functions for image super-resolution
Combining loss functions and GANs in an unsupervised way is useful for training image super-resolution models in the real world (Zhang et al., 2020). For instance, Zhang et al. used a novel loss function based image quality assessment in a GAN to obtain accurate texture information and more visual effects (Zhang et al., 2020). Besides, an encoder-decoder architecture is embedded in this GAN to mine more structure information for pursuing high-quality images of a generator from this GAN (Zhang et al., 2020). Han et al. depended on SAGAN and L1 loss in a GAN in an unsupervised manner to act multi-sequence structural MRI for detecting braining anomalies (Han et al., 2021a). Also, Zhang et al. fused a content loss into a GAN in an unsupervised manner to improve SR results of hyperspectral images (Zhang et al., 2021). Unsupervised GANs based prior knowledge and improved loss functions for image super-resolution can be summarized in Table 10.
Table 10. Unsupervised GANs based prior knowledge and improved loss functions for image super-resolution.
Models Methods Key words
DULGAN (Lin et al., 2018) GAN GAN with prior for image super-resolution
USROCTGAN (Das et al., 2020) GAN GAN with cycle consistency and identity mapping priors for image super-resolution
EIPGAN (Wang et al., 2021b) GAN GAN with remote sensing image prior for image super-resolution
URSGAN (Zhang et al., 2020) GAN GAN with prior based image quality assessment for remote sensing image super-resolution
MADGAN (Han et al., 2021a) SAGAN SAGAN with L1 loss for medical image super-resolution
DLGAN (Zhang et al., 2021) GAN GAN with a content loss for hyperspectral image super-resolution
4.3.4. Unsupervised GANs based multi-tasks for image super-resolution
Unsupervised GANs are good tools to address multi-tasks, i.e., noisy low-resolution image super-resolution. For instance, Prajapati et al. transferred a variational auto-encoder and the idea of quality assessment in a GAN to deal with image denoising and SR tasks (Prajapati et al., 2021). Cui et al. relied on low-pass-filter loss and weighted MR images in a GAN in an unsupervised GAN to mine texture information for removing noise and recovering resolution of MRI images (Cui et al., 2022). Cai et al. presented a pipeline that optimizes a periodic implicit GAN to obtain neural radiance fields for image synthesis and image super-resolution based on 3D (Cai et al., 2022). More unsupervised GANs based multi-tasks for image super-resolution can be presented in Table 11.
Table 11. Unsupervised GANs based multi-tasks for image super-resolution.
Models Methods Key words
VAEGAN (Prajapati et al., 2021) GAN GAN with variational auto-encoder and quality assessment idea for image super-resolution and denoising
ASLGAN (Cui et al., 2022) GAN GAN with low-pass-filter loss and weighted MR images for MRI image super-resolution and denoising
Pix2NeRF (Cai et al., 2022) GAN Optimizing a periodic implicit GAN for 3D-aware image synthesis and super-resolution
Pi-GAN (Chan et al., 2021a) GAN GAN with periodic activation functions for 3D-aware image synthesis and image super-resolution
5. Comparing performance of GANs for image super-resolution
To make readers conveniently know GANs in image super-resolution, we compare super-resolution performance of these GANs from datasets and experimental settings to quantitative and qualitative analysis in this section. More information can be shown as follows.
5.1. Datasets
Mentioned GANs can be divided into three kinds: supervised methods, semi-supervised methods and unsupervised methods for image super-resolution, which make datasets have three categories, training datasets and test datasets for supervised methods, semi-supervised methods and unsupervised methods. These datasets can be summarized as follows.
(1) Supervised GANs for image-resolution
Training datasets: CIFAR10 (Krizhevsky et al., 2009), STL (Wang and Tan, 2016), LSUN (Zhang et al., 2015), ImageNet (Russakovsky et al., 2015), Celeb A (Liu et al., 2015), DIV2K (Agustsson and Timofte, 2017), Flickr2K (Wang et al., 2019), OST (Wang et al., 2018a), CAT (Zhang et al., 2008) Market-1501 (Zheng et al., 2015), Duke MTMC-reID (Ristani et al., 2016; Zheng et al., 2017), GeoEye-1 satellite dataset (Ma et al., 2019), Whole slide images (WSIs) (Ma et al., 2020), MNIST (LeCun et al., 1998) and PASCAL2 (Everingham et al., 2010).
Test datasets: CIFAR10 (Krizhevsky et al., 2009), STL (Wang and Tan, 2016), LSUN (Zhang et al., 2015), Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), BSD100 (Martin et al., 2001), CELEBA (Liu et al., 2015), OST300 (Zhang et al., 2015), CAT (Zhang et al., 2008), PIRM datasets (Blau et al., 2018), Market-1501 (Zheng et al., 2015), GeoEye-1 satellite dataset (Ma et al., 2019), WSIs (Ma et al., 2020) MNIST (LeCun et al., 1998) and PASCAL2 (Everingham et al., 2010).
(2) Semi-supervised GANs for image-resolution
Training datasets: Market-1501 (Zheng et al., 2015), Tibia Dataset (Chen et al., 2018), Abdominal Dataset (McCollough et al., 2020), CUHK03 (Li et al., 2014), MSMT17 (Wei et al., 2018), LUNA (Setio et al., 2017), Data Science Bowl 2017 (DSB) (Kuan et al., 2017), UKDHP (Woodbridge et al., 2013), SG (Woodbridge et al., 2013) and UKBB (Bycroft et al., 2018).
Test datasets: Tibia Dataset (Chen et al., 2018), Abdominal Dataset (McCollough et al., 2020), CUHK03 (Li et al., 2014), Widerface (Yang et al., 2016a), LUNA (Setio et al., 2017), DSB (Kuan et al., 2017), SG (Woodbridge et al., 2013) and UKBB (Bycroft et al., 2018).
(3) Unsupervised GANs for image-resolution
Training datasets: CIFAR10 (Krizhevsky et al., 2009), ImageNet (Russakovsky et al., 2015), DIV2K (Agustsson and Timofte, 2017), DIV2K random kernel (DIV2KRK) (Agustsson and Timofte, 2017), Flickr2K (Wang et al., 2019), Widerface (Yang et al., 2016a), NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020), KADID-10K (Lin et al., 2019), DPED (Ignatov et al., 2017), DF2K (Wang et al., 2018b), NTIRE’ 2018 Blind-SR challenge (Timofte et al., 2018), LS3D-W (Bulat and Tzimiropoulos, 2017), CELEBA-HQ (Karras et al., 2017), LSUN-BEDROOM (Yu et al., 2015), ILSVRC2012 (Russakovsky et al., 2015; Nguyen et al., 2017), NTIRE 2020 (Lugmayr et al., 2020), 91-images (Yang et al., 2010), Berkeley segmentation (Martin et al., 2001), BSDS500 (Yang et al., 2016b), Training datasets of USROCTGAN (Fang et al., 2013, 2012), SD-OCT dataset (Fang et al., 2012), UC Merced dataset (Yang and Newsam, 2010), NWPU-RESIS45 (Cheng et al., 2017) and WHU-RS19 (Dai and Yang, 2010).
Test datasets: CIFAR10 (Krizhevsky et al., 2009), ImageNet (Russakovsky et al., 2015), Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), BSD100 (Martin et al., 2001), DIV2K (Agustsson and Timofte, 2017), DIV2KRK (Agustsson and Timofte, 2017), Urban100 (Huang et al., 2015), Widerface (Yang et al., 2016a), NTIRE 2020 (Lugmayr et al., 2020), NTIRE 2020 Real-world SR Challenge (Lugmayr et al., 2020), NTIRE 2020 Real World SR challenge validation dataset (Lugmayr et al., 2020), DPED (Ignatov et al., 2017), CELEBA-HQ (Karras et al., 2017), LSUN-BEDROOM (Yu et al., 2015), Test datasets of USROCTGAN (Fang et al., 2013, 2012), and Test datasets of USRGAN (Zhang et al., 2020).
These mentioned datasets about GANs for image super-resolution can be shown in Table 12. To make readers easier understand datasets of different methods via different GANs for different training ways on image super-resolutions, we conduct Table 13 to show their detailed information.
Table 12. Datasets (i.e., training datasets and test datasets) of GANs for image super-resolution.
Training ways Training datasets Test datasets
Supervised ways
CIFAR10 (Krizhevsky et al., 2009),
STL (Wang and Tan, 2016),
LSUN (Zhang et al., 2015),
ImageNet (Russakovsky et al., 2015),
Celeb A (Liu et al., 2015),
DIV2K (Agustsson and Timofte, 2017),
Flickr2K (Wang et al., 2019),
OST (Wang et al., 2018a),
CAT (Zhang et al., 2008),
Market-1501 (Zheng et al., 2015),
Duke MTMC-reID (Ristani et al., 2016; Zheng et al., 2017),
GeoEye-1 satellite dataset (Ma et al., 2019),
Whole slide images (WSIs) (Ma et al., 2020),
MNIST (LeCun et al., 1998),
PASCAL2 (Everingham et al., 2010),
Set5 (Bevilacqua et al., 2012),
Set14 (Zeyde et al., 2010),
BSD100 (Martin et al., 2001),
Urban100 (Huang et al., 2015).
CIFAR10 (Krizhevsky et al., 2009),
STL (Wang and Tan, 2016),
LSUN (Zhang et al., 2015),
Set5 (Bevilacqua et al., 2012),
Set14 (Zeyde et al., 2010),
BSD100 (Martin et al., 2001),
CELEBA (Liu et al., 2015),
CAT (Zhang et al., 2008),
OST300 (Wang et al., 2018a),
the PIRM datasets (Blau et al., 2018),
Market-1501 (Zheng et al., 2015),
GeoEye-1 satellite dataset (Ma et al., 2019),
WSIs (Ma et al., 2020),
MNIST (LeCun et al., 1998),
PASCAL2 (Everingham et al., 2010).
Semi-supervised ways
Tibia Dataset (Chen et al., 2018),
Abdominal Dataset (McCollough et al., 2020),
Market-1501 (Zheng et al., 2015),
CUHK03 (Li et al., 2014),
MSMT17 (Wei et al., 2018),
LUNA (Setio et al., 2017),
Data Science Bowl 2017 (DSB) (Kuan et al., 2017),
UKDHP (Woodbridge et al., 2013),
SG (Woodbridge et al., 2013),
UKBB (Bycroft et al., 2018).
Tibia Dataset (Chen et al., 2018),
Abdominal Dataset (McCollough et al., 2020),
Market-1501 (Zheng et al., 2015),
CUHK03 (Li et al., 2014)
LUNA (Setio et al., 2017),
DSB (Kuan et al., 2017),
SG (Woodbridge et al., 2013),
UKBB (Bycroft et al., 2018).
Unsupervised ways
DIV2K (Agustsson and Timofte, 2017),
Flickr2K (Wang et al., 2019),
Widerface (Yang et al., 2016a),
NTIRE-2020 Real-world SR Challenge validation dataset (Lugmayr et al., 2020),
KADID-10K (Lin et al., 2019),
DPED (Ignatov et al., 2017),
DF2K (Wang et al., 2018b),
NTIRE’2018 Blind-SR challenge (Timofte et al., 2018),
DIV2K random kernel (DIV2KRK) (Agustsson and Timofte, 2017),
LS3D-W (Bulat and Tzimiropoulos, 2017),
CIFAR10 (Krizhevsky et al., 2009),
ImageNet (Russakovsky et al., 2015),
CELEBA-HQ (Karras et al., 2017),
LSUN-BEDROOM (Yu et al., 2015),
NTIRE 2020 (Lugmayr et al., 2020),
91-images (Yang et al., 2010),
Berkeley segmentation (Martin et al., 2001), BSDS500 (Yang et al., 2016b),
Training datasets of USROCTGAN (Fang et al., 2013, 2012),
SD-OCT dataset (Fang et al., 2012),
UC Merced dataset (Yang and Newsam, 2010),
NWPU-RESIS45 (Cheng et al., 2017),
WHU-RS19 (Dai and Yang, 2010).
DIV2K (Agustsson and Timofte, 2017),
Set5 (Bevilacqua et al., 2012),
Set14 (Bevilacqua et al., 2012),
Urban100 (Huang et al., 2015),
BSD100 (Martin et al., 2001),
NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020),
NTIRE-2020 Real-world SR Challenge validation dataset (Lugmayr et al., 2020),
DPED (Ignatov et al., 2017),
DIV2KRK (Agustsson and Timofte, 2017),
Widerface (Yang et al., 2016a),
CIFAR10 (Krizhevsky et al., 2009),
ImageNet (Russakovsky et al., 2015),
CELEBA-HQ (Karras et al., 2017),
LSUN-BEDROOM (Yu et al., 2015),
NTIRE 2020 (Lugmayr et al., 2020),
Tests datasets of USROCTGAN (Fang et al., 2013),
Test datasets of USRGAN (Zhang et al., 2020).
Table 13. Different GANs on image super-resolution for different training ways.
Training ways Methods Training datasets Test datasets
Supervised ways
LAPGAN (Denton et al., 2015)
SRGAN (Ledig et al., 2017)
PGGAN (Karras et al., 2017)
ESRGAN (Wang et al., 2018b)
RaGAN (Jolicoeur-Martineau, 2018)
ESRGAN+ (Rakotonirina and Rasoanaivo, 2020)
SPGAN (Zhang and Ling, 2020)
SD-GAN (Ma et al., 2019)
PathSRGAN (Ma et al., 2020)
TGAN (Ding et al., 2019)
DGAN (Zareapoor et al., 2019)
G-GANISR (Shamsolmoali et al., 2019)
CIFAR10 (Krizhevsky et al., 2009), STL (Wang and Tan, 2016), LSUN (Zhang et al., 2015)
ImageNet (Russakovsky et al., 2015)
CIFAR10 (Krizhevsky et al., 2009), Celeb A (Liu et al., 2015)
DIV2K (Agustsson and Timofte, 2017), Flickr2K (Wang et al., 2019), OST (Wang et al., 2018a)
CIFAR10 (Krizhevsky et al., 2009), CAT (Zhang et al., 2008)
DIV2K (Agustsson and Timofte, 2017)
Market-1501 (Zheng et al., 2015), Duke MTMC-reID (Ristani et al., 2016; Zheng et al., 2017)
GeoEye-1 satellite dataset (Ma et al., 2019)
Whole slide images (WSIs) (Ma et al., 2020)
MNIST (LeCun et al., 1998), PASCAL2 (Everingham et al., 2010), CIFAR10 (Krizhevsky et al., 2009)
DIV2K (Agustsson and Timofte, 2017)
Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), BSD100 (Martin et al., 2001), Urban100 (Huang et al., 2015)
CIFAR10 (Krizhevsky et al., 2009), STL(Wang and Tan, 2016), LSUN (Zhang et al., 2015)
Set5 (Bevilacqua et al., 2012), Set14(Bevilacqua et al., 2012), BSD100 (Martin et al., 2001)
CELEBA (Liu et al., 2015), LSUN (Zhang et al., 2015), CIFAR10 (Krizhevsky et al., 2009)
Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), BSD100 (Martin et al., 2001), Urban100 (Huang et al., 2015)
CAT (Zhang et al., 2008)
BSD100 (Martin et al., 2001), Urban100 (Huang et al., 2015), OST300 (Wang et al., 2018a), Set5 (Bevilacqua et al., 2012), Set14 (Bevilacqua et al., 2012), the PIRM datasets (Blau et al., 2018)
Market-1501 (Zheng et al., 2015)
GeoEye-1 satellite dataset (Ma et al., 2019)
WSIs (Ma et al., 2020)
MNIST (LeCun et al., 1998), PASCAL2 (Everingham et al., 2010), CIFAR10 (Krizhevsky et al., 2009)
Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), CIFAR-10 (Krizhevsky et al., 2009), BSD100 (Martin et al., 2001)
Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), Urban100 (Huang et al., 2015)
Semi-supervised ways
GAN-CIRCLE (You et al., 2019a)
MSSR (Xia et al., 2021b)
CTGAN (Jiang et al., 2020)
Gemini-GAN (Savioli et al., 2021)
Tibia Dataset (Chen et al., 2018), Abdominal Dataset (McCollough et al., 2020)
Market-1501 (Zheng et al., 2015), CUHK03 (Li et al., 2014), MSMT17 (Wei et al., 2018)
LUNA (Setio et al., 2017), Data Science Bowl 2017 (DSB) (Kuan et al., 2017)
UKDHP (Woodbridge et al., 2013), SG (Woodbridge et al., 2013), UKBB (Bycroft et al., 2018)
Tibia Dataset (Chen et al., 2018), Abdominal Dataset (McCollough et al., 2020)
Market-1501 (Zheng et al., 2015), CUHK03 (Li et al., 2014)
LUNA (Setio et al., 2017), DSB (Kuan et al., 2017)
SG (Woodbridge et al., 2013), UKBB (Bycroft et al., 2018)
Unsupervised ways
CinCGAN (Yuan et al., 2018)
DNSR (Zhao et al., 2018)
MCinCGAN (Zhang et al., 2019c)
RWSR-CycleGAN (Kim et al., 2020)
USISResNet (Prajapati et al., 2020)
ULRWSR (Lugmayr et al., 2019)
KernelGAN (Bell-Kligler et al., 2019)
FG-SRGAN (Lian et al., 2019)
TrGAN (Wang et al., 2020b)
PDCGAN (Miyato and Koyama, 2018)
CycleSR (Chen et al., 2020)
DULGAN (Lin et al., 2018)
USROCTGAN (Das et al., 2020)
URSGAN (Zhang et al., 2020)
DIV2K (Agustsson and Timofte, 2017)
DIV2K (Agustsson and Timofte, 2017), Flickr2K (Wang et al., 2019), Widerface (Yang et al., 2016a)
DIV2K (Agustsson and Timofte, 2017)
NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020)
NTIRE-2020 Real-world SR Challenge validation dataset (Lugmayr et al., 2020), DIV2K (Agustsson and Timofte, 2017), Flickr2k (Wang et al., 2019), KADID-10K (Lin et al., 2019)
DPED (Ignatov et al., 2017), DF2K (Wang et al., 2018b), DIV2K (Agustsson and Timofte, 2017), Flickr2K (Wang et al., 2019)
NTIRE’2018 Blind-SR challenge (Timofte et al., 2018), DIV2K random kernel (DIV2KRK) (Agustsson and Timofte, 2017)
LS3D-W (Bulat and Tzimiropoulos, 2017)
CIFAR10 (Krizhevsky et al., 2009), ImageNet (Russakovsky et al., 2015), CELEBA-HQ (Karras et al., 2017), LSUN-BEDROOM (Yu et al., 2015)
ImageNet (Russakovsky et al., 2015), ILSVRC2012 (Russakovsky et al., 2015; Nguyen et al., 2017), CIFAR10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009)
DIV2K (Agustsson and Timofte, 2017), NTIRE 2020 (Lugmayr et al., 2020)
91-images (Yang et al., 2010), Berkeley segmentation (Martin et al., 2001), BSDS500 (Yang et al., 2016b)
Training datasets of USROCTGAN (Fang et al., 2013, 2012), SD-OCT dataset (Fang et al., 2012)
UC Merced dataset (Yang and Newsam, 2010), NWPU-RESIS45 (Cheng et al., 2017), WHU-RS19 (Dai and Yang, 2010)
DIV2K (Agustsson and Timofte, 2017)
Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), Urban100 (Huang et al., 2015), BSD100 (Martin et al., 2001), DIV2K (Agustsson and Timofte, 2017)
DIV2K (Agustsson and Timofte, 2017)
NTIRE 2020 Real World SR challenge (Lugmayr et al., 2020)
NTIRE-2020 Real-world SR Challenge validation dataset (Lugmayr et al., 2020)
DPED (Ignatov et al., 2017), DIV2K (Agustsson and Timofte, 2017)
DIV2KRK (Agustsson and Timofte, 2017)
Widerface (Yang et al., 2016a)
CIFAR10 (Krizhevsky et al., 2009), ImageNet (Russakovsky et al., 2015), CELEBA-HQ (Karras et al., 2017), LSUN-BEDROOM (Yu et al., 2015)
ImageNet (Russakovsky et al., 2015)
DIV2K (Agustsson and Timofte, 2017), NTIRE 2020 (Lugmayr et al., 2020)
Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010)
Test datasets of USROCTGAN (Fang et al., 2013)
Test datasets of USRGAN (Zhang et al., 2020)
5.2. Environment configurations
In this section, we compare the differences of environment configurations between different GANs via different training ways (i.e., supervised, semi-supervised and unsupervised ways) for image super-resolution, which contain batch size, scaling factors, deep learning framework, learning rate and iteration. That can make readers easier to conduct experiments with GANs for image super-resolution. Their information can be listed as shown in Table 14 as follows.
Table 14. Environment configurations of different GANs for image super-resolution.
Training ways Methods Batchsize Scaling factors Framework Learning rates Iteration
Supervised ways
LAPGAN (Denton et al., 2015)
ESRGAN (Wang et al., 2018b)
ESRGAN+ (Rakotonirina and Rasoanaivo, 2020)
SPGAN (Zhang and Ling, 2020)
SD-GAN (Ma et al., 2019)
TGAN (Ding et al., 2019)
DGAN (Zareapoor et al., 2019)
G-GANISR (Shamsolmoali et al., 2019)
GMGAN (Zhu et al., 2020)
16
16
16
16
9
32
16
×2
×4
×4
×2, ×4, ×8 and ×16
×3
×2
×4, ×6 and ×8
×4, ×6 and ×8
×4
PyTorch (Paszke et al., 2019)
PyTorch (Paszke et al., 2019)
PyTorch (Paszke et al., 2019)
TensorFlow (Goldsborough, 2016)
TensorFlow (Goldsborough, 2016)
TensorFlow (Goldsborough, 2016)
TensorFlow (Goldsborough, 2016)
PyTorch (Paszke et al., 2019) and TensorFlow (Goldsborough, 2016)
PyTorch (Paszke et al., 2019)
0.02
$10^{-4}$, $10^{-4}$
$10^{-4}$
$10^{-3}$
$10^{-3}$
$10^{-4}$
0.125
$10^{-4}$
$10^{-4}$
200K, 300K
300K
2K
500
25K
800K
100K
Semi-supervised ways
MSSR (Xia et al., 2021b)
PSSR (Maeda, 2020)
CTGAN (Jiang et al., 2020)
64
16
16
×4
×4
×4
PyTorch (Paszke et al., 2019)
PyTorch (Paszke et al., 2019)
TensorFlow (Goldsborough, 2016)
$10^{-3}$, 0.01
$10^{-4}$
$10^{-4}$
300K
60K
Unsupervised ways
CinCGAN (Yuan et al., 2018)
DNSR (Zhao et al., 2018)
MCinCGAN (Zhang et al., 2019c)
RWSR-CycleGAN (Kim et al., 2020)
USISResNet (Prajapati et al., 2020)
ULRWSR (Lugmayr et al., 2019)
KernelGAN (Bell-Kligler et al., 2019)
InGAN (Shocher et al., 2018)
PETSRGAN (Mahapatra et al., 2017)
PDCGAN (Miyato and Koyama, 2018)
USROCTGAN (Das et al., 2020)
URSGAN (Zhang et al., 2020)
16
16
1
16
32
1
10, 20
64
64
×4
×2, ×4
×2, ×4 and ×8
×4
×4
×4
×2, ×4
×2
×4
×4
×4
×2, ×4
PyTorch (Paszke et al., 2019)
TensorFlow (Goldsborough, 2016)
TensorFlow (Goldsborough, 2016)
PyTorch (Paszke et al., 2019)
TensorFlow (Goldsborough, 2016)
PyTorch (Paszke et al., 2019)
PyTorch (Paszke et al., 2019)
TensorFlow (Goldsborough, 2016)
TensorFlow (Goldsborough, 2016)
PyTorch (Paszke et al., 2019)
$10^{-4}$
$10^{-4}$
$10^{-4}$
$10^{-4}$
$10^{-4}$
$10^{-4}$, 2×$10^{-4}$
$10^{-4}$
$10^{-4}$
$10^{-4}$
$10^{-4}$
$10^{-4}$
$10^{-4}$
400K
106
500K
300K
120K
50K
3K
20K
850K
68K
5.3. Experimental results
To make readers understand the performance of different GANs on image super-resolution, we use quantitative analysis and qualitative analysis to evaluate super-resolution effects of these GANs. Quantitative analysis is PSNR and SSIM of different methods via three training ways on different datasets for image super-resolution, running time and complexities of different GANs on image super-resolution. Qualitative analysis is used to evaluate qualities of recovered images.
5.3.1. Quantitative analysis of different GANs for image super-resolution
We use SRGAN (Ledig et al., 2017), PGGAN (Karras et al., 2017), ESRGAN (Wang et al., 2018b), ESRGAN+ (Rakotonirina and Rasoanaivo, 2020), DGAN (Zareapoor et al., 2019), G-GANISR (Shamsolmoali et al., 2019), GMGAN (Zhu et al., 2020), SRPGAN (Ledig et al., 2017), DNSR (Zhao et al., 2018), DULGAN (Lin et al., 2018), CinCGAN (Yuan et al., 2018),MCinCGAN (Zhang et al., 2019c), USISResNet(Prajapati et al., 2020), ULRWSR (Lugmayr et al., 2019), KernelGAN (Bell-Kligler et al., 2019) and CycleSR (Chen et al., 2020) in one training way from supervised, semi-supervised and unsupervised ways on a public dataset from Set14 (Zeyde et al., 2010), BSD100 (Martin et al., 2001) and DIV2K (Agustsson and Timofte, 2017) to test performance for different scales in image super-resolution as shown in Table 15. For instance, ESRGAN (Wang et al., 2018b) outperforms SRGAN (Ledig et al., 2017) in terms of PSNR and SSIM in a supervised ways on Set 14 for ×2, which shows that ESRGAN has obtained better super-resolution performance for ×2. More information of these GANs can be shown in Table 15.
Table 15. PSNR and SSIM of different GANs via different training ways on Set14, BSD100 and DIV2K for image super-resolution.
Training ways Methods Datasets Scale PSNR SSIM
Supervised ways SRGAN (Ledig et al., 2017) Set14 (Zeyde et al., 2010) ×2 32.14 0.8860
×4 26.02 0.7379
PGGAN (Karras et al., 2017) ×6 29.54 0.8301
×8 28.14 0.8094
ESRGAN (Wang et al., 2018b) ×2 33.62 0.9150
×4 30.50 0.7620
ESRGAN+ (Rakotonirina and Rasoanaivo, 2020) ×4 19.79 -
DGAN (Zareapoor et al., 2019) ×4 31.62 0.9166
×6 28.62 0.9003
×8 26.85 0.8911
G-GANISR (Shamsolmoali et al., 2019) ×4 29.67 -
×6 30.56 0.8881
×8 28.07 0.8803
GMGAN (Zhu et al., 2020) ×4 26.37 0.7055
SRPGAN (Ledig et al., 2017) ×6 26.19 0.8187
×8 29.17 0.8733
SRGAN (Ledig et al., 2017) BSD100 (Martin et al., 2001) ×2 31.89 0.8760
×4 25.16 0.6688
ESRGAN (Wang et al., 2018b) ×2 31.99 0.8870
×4 27.69 0.7120
DGAN (Zareapoor et al., 2019) ×4 31.53 0.9105
×6 29.62 0.8937
×8 27.85 0.8811
G-GANISR (Shamsolmoali et al., 2019) ×4 28.12 -
×6 31.23 0.9273
×8 29.18 0.9065
GMGAN (Zhu et al., 2020) ×4 25.46 0.6592
SRPGAN (Ledig et al., 2017) ×6 27.48 0.8652
×8 23.18 0.8504
SRGAN (Ledig et al., 2017) DIV2K (Agustsson and Timofte, 2017) ×2 25.08 0.7007
×4 28.09 0.8210
ESRGAN (Wang et al., 2018b) ×4 28.68 0.8530
Semi-supervised ways PSSR (Maeda, 2020) DIV2K (Agustsson and Timofte, 2017) ×4 21.32 0.5541
Unsupervised ways DNSR (Zhao et al., 2018) Set14 (Zeyde et al., 2010) ×2 33.83 0.9220
×4 31.76 0.8910
DULGAN (Lin et al., 2018) ×4 27.39 0.7412
DNSR (Zhao et al., 2018) BSD100 (Martin et al., 2001) ×2 32.24 0.9010
×4 25.69 0.7880
CinCGAN (Zhang et al., 2019c) DIV2K (Agustsson and Timofte, 2017) ×2 25.61 0.6957
×4 25.32 0.6705
×8 24.58 0.6581
DNSR (Zhao et al., 2018) ×4 28.87 0.8650
MCinCGAN (Zhang et al., 2019c) ×2 26.15 0.7020
×4 25.51 0.6878
×8 24.79 0.6618
USISResNet (Prajapati et al., 2020) ×4 21.22 0.5760
ULRWSR (Lugmayr et al., 2019) ×4 23.30 0.6200
KernelGAN (Bell-Kligler et al., 2019) ×2 30.363 0.8669
×4 26.810 0.7316
CycleSR (Chen et al., 2020) ×4 23.807 0.5930
Running time and complexity are important indexes to evaluate performance of image super-resolution techniques in the real devices (Tian et al., 2020). According to that, we conduct experiments of four GANs (i.e., ESRGAN (Wang et al., 2018b), PathSRGAN (Ma et al., 2020), RankSRGAN (Zhang et al., 2019b) and KernelGAN (Bell-Kligler et al., 2019)) on two low-resolution images with sizes and for ×4 to test running time and compute parameters of different GANs. The conducted experiments have the following experimental environments. They can run on Ubuntu of 20.04.1, CPU of AMD EPYC ROME 7502P with 32 cores and Memory of 128G via PyTorch of 1.10.1 (Paszke et al., 2019). Besides, they depend on a NVIDIA GeFore RTX 3090 with cuda of 11.1 and cuDNN of. 8.0.4. In Table 16, we can see that ESRGAN (Wang et al., 2018b) has slower speed than that of PathSRGAN for ×4 on image super-resolution. However, it uses less parameters than that of PathSRGAN for ×4 on image super-resolution. Thus, ESRGAN is competitive with PathSRGAN for image super-resolution. More information of different GANs for image super-resolution in terms of running time and parameters can be shown in Table 16.
Table 16. Running time and parameters of different GANs for ×4.
Methods Testing time (s) Parameters
ESRGAN (Wang et al., 2018b) 9.7607 1.670×$10^{7}$
PathSRGAN (Ma et al., 2020) 9.1608 2.433×$10^{7}$
RankSRGAN (Zhang et al., 2019b) 7.3818 1.554×$10^{6}$
KernelGAN (Bell-Kligler et al., 2019) 251.00 1.816×$10^{5}$
5.3.2. Qualitative analysis of different GANs for image super-resolution
To test visual effects of different GANs for image super-resolution, we choose Bicubic, ESRGAN (Wang et al., 2018b), RankSRGAN (Zhang et al., 2019b), KernelGAN (Bell-Kligler et al., 2019) and PathSRGAN (Ma et al., 2020) to conduct experiments to obtain high-quality images for ×4. To further observe these images, we choose an area of predicted images from these GANs to amplify it as an observation area. Observation area is clearer, corresponding method has good superior SR performance. For example, ESRGAN (Wang et al., 2018b) is clearer than that of PathSRGAN (Ma et al., 2020) on an image from the BSD100 in Fig. 9 and Set14 in Fig. 10 for ×4, which show that the ESRGAN is more effective in image super-resolution.
Figure 9. Visual images of different GANs on an image of BSD100 for ×4: (a) original image, (b) Bicubic, (c) ESRGAN, (d) RankSRGAN, (e) KernelGAN, and (f) PathSRGAN.
[…]
Qualitative experiment on BSD100.
Figure 10. Visual images of different GANs on an image of Set14 for ×4: (a) original image, (b) Bicubic, (c) ESRGAN, (d) RankSRGAN, (e) KernelGAN, and (f) PathSRGAN.
[…]
Qualitative experiment on Set14.
6. Challenges and directions of GANs for image super-resolution
Variations of GANs have achieved excellent performance in image super-resolution. Accordingly, we provide an overview of GANs for image super-resolution to offer a guide for readers to understand these methods. In this section, we analyze challenges of current GANs for image super-resolution and give corresponding solutions to facilitate the development of GANs for image super-resolution.
Although GANs perform well in image super-resolution, they suffer from the following challenges.
1) Unstable training. Due to the confrontation between generator and discriminator, GANs are unstable in the training process.
2) Large computational resources and high memory consumption. A GAN is composed of a generator and discriminator, which may increase computational costs and memory consumption. This may lead to a higher demand on digital devices.
3) High-quality images without references. Most of existing GANs relied on paired high-quality images and low-resolution images to train image super-resolution models, which may be limited by digital devices in the real world.
4) Complex image super-resolution. Most of GANs can deal with a single task, i.e., image super-resolution and synthetic noisy image super-resolution, etc. However, collected images by digital cameras in the real world suffer from drawbacks, i.e., low-resolution and dark-lighting images, complex noisy and low-resolution images. Besides, digital cameras have higher requirement on the combination of image low-resolution and image recognition. Thus, existing GANs for image super-resolution cannot effectively repair low-resolution images of mentioned conditions.
5) Metrics of GANs for image super-resolution. Most of existing GANs used PSNR and SSIM to test super-resolution performance of GANs. However, PSNR and SSIM cannot fully measure restored images. Thus, finding effective metrics is very essential about GANs for image super-resolution.
To address these problems, some potential research points about GANs for image super-resolution are stated below.
1) Enhancing a generator and discriminator extracts salient features to enhance stabilities of GANs on image super-resolution. For example, using attention mechanism (i.e., Transformer (Han et al., 2021b)), residual learning operations, concatenation operations act a generator and discriminator to extract more effective features to enhance stabilities for accelerating GAN models in image super-resolution.
2) Designing lightweight GANs for image super-resolution. Reducing convolutional kernels, group convolutions, the combination of prior and shallow network architectures can decrease the complexities of GANs for image super-resolution.
3) Using self-supervised methods can obtain high-quality reference images.
4) Combining attributes of different low-level tasks, decomposing complex low-level tasks into a single low-level task via different stages in different GANs repairs complex low-resolution images, which can help high-level vision tasks.
5) Using image quality assessment techniques as metrics evaluates quality of precited images from different GANs.
7. Conclusion
In this paper, we analyze and summarize GANs for image super-resolution. First, we introduce developments of GANs. Then, we present GANs in big samples and small samples for image applications, which can make readers easier GANs. Next, we give differences of GANs based optimization methods and discriminative learning for image super-resolution in terms of supervised, semi-supervised and unsupervised manners. Subsequently, we compare the performance of these popular GANs on public datasets via quantitative and qualitative analysis in SISR. Finally, we highlight challenges of GANs and potential research points on SISR.
Acknowledgements.
This work was supported in part by National Natural Science Foundation of China under Grant 62201468, in part by the China Postdoctoral Science Foundation Grant 2022TQ0259, in part by the Jiangsu Provincial Double–Innovation Doctor Program Grant JSSCBC20220942. The authors appreciate informative and insightful comments provided by anonymous editors and reviewers, which significantly improve the quality of this survey.